James is twice as old as john,if 16 is added to john's age and 16 is subtracted from james age,their ages will then be equal.what are their present ages?
Answer by sachi(548)

let present age of john is x
then james is 2x
as per question
16+x=2x-16
or 2x-x=16+16
or x=32= john's age
& james is 64 yrs ans